OpenStudy (anonymous):

Simplify this rational expression. 12x+36/48x

mathslover (mathslover):

did u mean this : \[\huge{12x+\frac{36x}{48}}\]

OpenStudy (anonymous):

No I mean \[12\chi+38\div48x\] The 12x+38 is the top part of the fraction then it's over 48x

OpenStudy (anonymous):

 Like this.

mathslover (mathslover):

ok see when u r given that : \(\huge{\frac{a+b}{c}}\) then it can be written as\(\huge{\frac{a}{c}+\frac{b}{c}}\) similarly here also : given : \(\huge{\frac{12x+36}{48x}}\) u can write this as : \(\huge{\frac{12x}{48x}+\frac{36}{48x}}\)

OpenStudy (anonymous):

OKay so.. would the answer be..... (x+3)/(4x)?

mathslover (mathslover):

since it is given to simplify i will say that u can write this as : \[\huge{\frac{1}{4}+\frac{3}{4x}}\] or that one u said ... both r right but ur one is more correct

OpenStudy (anonymous):

Yay(: Thanks!